kindly help me with this other problem. PLEASE
A father is 20 years older than his son. In 4 years time he will be twice as old as his son. Determine their present ages.

Answer by ankor@dixie-net.com(15622)

write an equation for each statement:
:A father is 20 years older than his son."
f = s + 20
:
" In 4 years time he will be twice as old as his son."
f + 4 = 2(s+4)
f = 4 = 2s + 8
f = 2s + 8 - 4
f = 2s +4
:
Determine their present ages.
:
Substitute (s+20) for f in the above equation, (from the 1st equation); find s:
s + 20 = 2s + 4
20 - 4 = 2s - s
16 = s
:
The son is 16 and the father is (16+20) = 36
:
:
Check solution in the statement:
"In 4 years time he will be twice as old as his son. "
36 + 4 = 2(16 + 4); confirms our solution